# Pólya’s conjecture up to $\epsilon$-loss and quantitative estimates for the remainder of Weyl’s law 00footnotetext: 2020 Mathematics Subject Classification. Primary 35P15; Secondary 35P20, 42B37. Key words and phrases: Dirichlet eigenvalue, Weyl law, Pólya conjecture

### v2

@@
where $\Lambda(\epsilon,\Omega)$ is given explicitly. This estimate is based on a uniform estimate on the remainder of the Weyl law with explicit constants so that it reduces the $\epsilon$-loss version of Pólya’s conjecture to a computational problem. Our arguments in deriving such uniform estimates yield also, in all dimensions $n\geq 2$, classes of domains that may even have rather irregular shapes or boundaries but satisfy Pólya’s conjecture.
@@ 1.1 Weyl’s law and Pólya’s conjecture
(1.1) $\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+o(\lambda^{n/2}),$
@@
(1.2) $R_{\Omega}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)-\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$
@@
(1.3) $\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}-\frac{|\partial\Omega|}{2^{n+1}\pi^{\frac{d-1}{2}}\Gamma(\frac{n+1}{2})}\lambda^{\frac{n-1}{2}}+o(\lambda^{\frac{n-1}{2}}),$
@@
However, the above mentioned results for the remainder $R_{\Omega}(\lambda)$ state that certain asymptotic estimates valid only for large $\lambda$’s without quantifying the sizes of such $\lambda$’s, and consequently they do not give quantitative estimate on eigenvalues. [^1] After the first version of the present article was posted on arXiv, R. Frank and S. Larson has kindly informed us recent works [9] and [12], in which several uniform estimates for the remainder were already established, see the added remark at the end of the introduction.
@@
(1.4) $\frac{|\Omega|^{2/n}\omega(n)^{2/n}}{(2\pi)^{2}}\sum_{i=1}^{k}\lambda_{i}\geq\frac{n}{n+2}k^{\frac{n+2}{n}},$
@@
(1.5) $\left(\frac{n+2}{n}\right)^{n/2}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}^{n/2}\geq k,$
@@
Our main purpose of this paper is to reduce the $\epsilon$-loss version of Pólya’s conjecture to a computational problem, and to give a solution to Question 1.1 with explicit constants, on Lipschitz domains. Moreover, we shall also provide an approach to the Pólya conjecture that might be applied to general smooth domains, and provide in all dimensions, $n\geq 2$, a class of domains that satisfy Pólya’s conjecture.
@@ 1.2 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
(1.6) |$\left|\{x\in{{{\mathbb{R}}}^{n}}:\,{\mathop{\mathrm{missing}}{\,dist\,}}(x,\partial\Omega)\leq\epsilon\}\right|\leq C_{Lip}(\Omega)\epsilon|\partial\Omega|.$
@@
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded Lipschitz domain, $n\geq 2$. For any $10<\epsilon<$ 1, it holds for all $\lambda_{k}(\Omega)\geq\Lambda(\epsilon,\Omega)$ that
@@
(1.7) $\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}C_{Lip}(\Omega)\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$
@@
(1.8) $\frac{2n\mathrm{diam}(\Omega)^{n-1}}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$
@@
Theorem 1.3 will follow from our following uniform estimate for the remainder of the Weyl’s law. Let us start from the following definition.
@@
We have the following uniform estimate on Lipschitz domains. In what follows, for a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$.
@@
(1.9) $C(n,\Omega,\lambda_{k}(\Omega))=C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\overline{\Omega})|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|.$
@@
(1.10) $k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
@@
(1.11) $C_{1}(n)=\begin{cases}\frac{2}{3}&n=1,\\ \frac{3\pi}{16}&n=2,\\ \frac{\pi\mathscr{B}(\frac{n}{2}-1,2)}{6\mathscr{B}(\frac{n}{2}-1,5/2)}&n=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n}{2}-1,2)}{16\mathscr{B}(\frac{n}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
Obviously, both quantities $\frac{r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}$ and $\frac{r_{in}(\Omega)}{|I_{min}|}$ are not larger than $1/2$, and therefore can be replaced by $1/2$. Moreover, note that if $|\partial(\mathscr{R}\setminus\overline{\Omega})|$ is small compared to $|\mathscr{R}_{n-1}|$, we then have that Pólya’s conjecture holds for the first some eigenvalues. Our quantitative estimate is consistent to the asymptotic estimates established by Courant [5] and Courant-Hilbert [6].
@@ 1.3 An approach to Pólya’s conjecture and concrete examples
By a rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I\subset{{{\mathbb{R}}}^{n}}$, $n\geq 2$, we mean $\mathscr{R}_{n-1}\subset{\mathbb{R}}^{n-1}$ is a rectangle, and $I\subset{\mathbb{R}}$ is an open interval.
@@
Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$. Assume that $\Omega$ is a piecewise smooth domain contained in $\mathscr{R}$. There exists $C_{P}=C(\Omega,n)$ such that if $|\mathscr{R}_{n-1}|\geq C_{P}$ then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\Omega}$.
@@
Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$, that is contained in $\mathscr{R}$. If $|\mathscr{R}_{n-1}|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where
(1.12) $C_{2}(n)=\begin{cases}2\sqrt{2}\pi,&n=1,\\ 9\sqrt{3},&n=2,\\ \frac{6(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{\mathscr{B}(\frac{n}{2}-1,2)}&n=3,4,\\ \frac{16(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{3\mathscr{B}(\frac{n}{2}-1,2)}\left(\frac{256}{247}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
For $M\geq 2$, by $\tilde{\Omega}$ $M$-tiles $\Omega$ we mean there are $M$ isometries of $\tilde{\Omega}$, $\{\tilde{\Omega}_{k}\}_{1\leq k\leq M}$, with non-overlapped interior, such that $\cup_{1\leq k\leq M}\tilde{\Omega}_{k}\subset\Omega$ and $|\Omega|=M|\tilde{\Omega}|$.
@@
Let $M\geq 2$. Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$, and $\Omega$ $M$-tiles $\mathscr{R}$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$, that is contained in $\Omega$. If $|\mathscr{R}_{n-1}|\geq C_{2}(n-1)M\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$.
Note that $M$ isometries of the domain $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$ corresponds to the case that removing $M$ copies of $\{Q_{k}\}_{k\in{\mathbb{N}}}$ from $\mathscr{R}$, the conclusion is obvious.
@@
(i) Let us take $\ell(k)=2^{\lfloor k/2\rfloor}$ for $k\geq 2$, $\ell(1)$=1, and $\{Q_{1,1},\,\{Q_{k,j}\}_{k\geq 2,1\leq j\leq\ell(k)}\}$ be a sequence of cubes, where $\ell(Q_{k,j})=2^{1-k}$. Then $\mathscr{S}_{Q}\leq 2(2+\sqrt{2})$. So for a rectangle $\Omega_{1}\times\Omega_{2}$, $|\Omega_{2}|>1$ and $|\Omega_{1}|>8(1+\sqrt{2})\pi$, one can arbitrarily remove any domain, that is composed by all or some of these cubes $\{Q_{1,1},\,\{Q_{k,j}\}_{k\geq 2,1\leq j\leq\ell(k)}\}$ with non-overlapped interior, from $\Omega_{1}\times\Omega_{2}$ such that the remaining set satisfies Pólya’s conjecture. See Figure 1 for some examples.
(a) Rectangle with a diamond and a room-corridor removed

### v3

@@
where $\Lambda(\epsilon,\Omega)$ is given explicitly. This reduces the $\epsilon$-loss version of Pólya’s conjecture to a computational problem. This estimate is based on quantitative estimates on the remainder of the Weyl law with explicit constants, which we give a new proof without using Neumann eigenvalues. Our arguments in deriving such uniform estimates yield also, in all dimensions $n\geq 2$, classes of domains that may even have rather irregular shapes or boundaries but satisfy Pólya’s conjecture. Another key observation is that on strip-tiling domains (and therefore any triangles for instance) one actually has better eigenvalue estimates than Pólya conjectured.
@@ 1.1 Weyl’s law and Pólya’s conjecture
$\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+o(\lambda^{n/2}),$ (1.1)
@@
$R_{\Omega}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)-\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$ (1.2)
@@
$\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}-\frac{|\partial\Omega|}{2^{n+1}\pi^{\frac{d-1}{2}}\Gamma(\frac{n+1}{2})}\lambda^{\frac{n-1}{2}}+o(\lambda^{\frac{n-1}{2}}),$ (1.3)
@@
However, the above mentioned results for the remainder $R_{\Omega}(\lambda)$ state that certain asymptotic estimates valid only for large $\lambda$’s without quantifying the sizes of such $\lambda$’s, and consequently they do not give quantitative estimate on eigenvalues.
@@
It was proved earlier by Netrusov and Safarov [32], that a quantitative bound up to an extra logarithm term of the remainder of Weyl’s law holds on Lipshcitz type domains, see also [9, 12].  [^1] After the first version of the present article was posted on arXiv, R. Frank and S. Larson has kindly informed us recent works [9] and [12], X. He informed us the work [32], in which several uniform estimates for the remainder were already established.
@@
$\frac{|\Omega|^{2/n}\omega(n)^{2/n}}{(2\pi)^{2}}\sum_{i=1}^{k}\lambda_{i}\geq\frac{n}{n+2}k^{\frac{n+2}{n}},$ (1.4)
@@
$\left(\frac{n+2}{n}\right)^{n/2}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}^{n/2}\geq k,$ (1.5)
@@
Our main purpose of this paper is to establish an approach to Pólya’s conjecture on more general domains (than tiling domains and products). More precisely, given a Lipschitz domain $\Omega$, we can choose a larger domain $T$, so that by using the monotone property of Dirichlet eigenvalues, one has
$\mathcal{N}^{D}_{\Omega}(\lambda)\leq\mathcal{N}^{D}_{T}(\lambda)-\mathcal{N}^{D}_{T\setminus\overline{\Omega}}(\lambda).$
The Pólya conjecture will follow if one can prove, on one side $T$ satisfies better eigenvalue estimates than Pólya conjectured, and one the other side, $\mathcal{N}^{D}_{T\setminus\overline{\Omega}}(\lambda)$ has a reasonable good lower bound.
On one side, we shall prove that strip-tiling domains and certain product domains satisfy better eigenvalue estimates than Pólya conjectured. On the other side, we shall prove on certain domains composed by cubes of different scales the eigenvalue counting functions do have a good lower bound. These enable us to provide in all dimensions, $n\geq 2$, a class of domains that satisfy Pólya’s conjecture.
For general Lipschitz domains, we did not have a good control of lower bound of $\mathcal{N}^{D}_{T\setminus\overline{\Omega}}(\lambda)$, and our method only yields a new proof of quantitative Courant type estimate for the remainder. Let us emphasize that quantitative Courant type estimate for the remainder has already obtained in [32] (see also [7]) via Courant’s method of Dirichlet-Neumann bracketing, our estimate (see Theorem 1.5 and Corollary 1.6 below) give slightly better constants.
If one can further obtain better estimate on the lower bound of $\mathcal{N}^{D}_{T\setminus\overline{\Omega}}(\lambda)$, then one can combine the refined estimate on $\mathcal{N}^{D}_{T}(\lambda)$ (see Remark 2.5 below, which behaves better as $\lambda\to\infty$), to reduce Pólya’s conjecture to a computational problem. In this paper, we can only reduce the $\epsilon$-loss version of Pólya’s conjecture to a computational problem.
@@ 1.2 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
$\left|\{x\in{{{\mathbb{R}}}^{n}}:\,{\mathop{\mathrm{\,dist\,}}}(x,\partial\Omega)\leq\epsilon\}\right|\leq C_{Lip}(\Omega)\epsilon|\partial\Omega|.$ (1.6)
@@
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded Lipschitz domain, $n\geq 2$. For any $0<\epsilon<1$, it holds for all $\lambda_{k}(\Omega)\geq\Lambda(\epsilon,\Omega)$ that
@@
$\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}C_{Lip}(\Omega)\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$ (1.7)
@@
$\frac{2n\mathrm{diam}(\Omega)^{n-1}}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$ (1.8)
@@
Theorem 1.3 follows directly from our following uniform estimate for the remainder of the Weyl’s law. Let us start from the following definition.
@@
We have the following uniform estimate on Lipschitz domains. Let us emphasize once more the following Courant type quantitative estimate has been obtained already in [32, Theorem 1.8] via Courant’s method of Dirichlet-Neumann bracketing. Our method for the proof of Theorem 1.5 and Corollary 1.6 below does not use Neumann eigenvalue and give better constant estimates.
In what follows, for a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$.
@@
$C(n,\Omega,\lambda_{k}(\Omega))=C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\overline{\Omega})|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|.$ (1.9)
@@
$k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$ (1.10)
@@
$C_{1}(n)=\begin{cases}\frac{2}{3}&n=1,\\ \frac{3\pi}{16}&n=2,\\ \frac{\pi\mathscr{B}(\frac{n}{2}-1,2)}{6\mathscr{B}(\frac{n}{2}-1,5/2)}&n=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n}{2}-1,2)}{16\mathscr{B}(\frac{n}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$ (1.11)
Obviously, both quantities $\frac{r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}$ and $\frac{r_{in}(\Omega)}{|I_{min}|}$ are not larger than $1/2$, and therefore can be replaced by $1/2$. Moreover, note that if $|\partial(\mathscr{R}\setminus\overline{\Omega})|$ is small compared to $|\mathscr{R}_{n-1}|$, we then have that Pólya’s conjecture holds for the first some eigenvalues.
@@ 1.3 An approach to Pólya’s conjecture and concrete examples
Definition 1.7 (Strip-tiling domains).
Let $\Omega$ be a bounded Lipschitz domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$. Up to rotations and translations, we assume that $\Omega$ is contained in two hyperplanes $\mathbb{R}^{n-1}\times\{0\}$ and $\mathbb{R}^{n-1}\times\{w(\Omega)\}$ for some $w(\Omega)\in(0,\infty)$. We say that $\Omega$ is a strip-tiling domain, if one can use isometries of $\Omega$, obtained by rotating $\Omega$ only in the first n1n-1 directions, to cover $\mathbb{R}^{n-1}\times(0,w(\Omega))$ without overlapped interiors.
Obviously, a rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I\subset{{{\mathbb{R}}}^{n}}$, $n\geq 2$ is a strip-tiling domain. The following Figure 1 provide more examples of general strip-tiling domains.
Figure 1: Examples of strip-tiling domains in two dimension
Remark 1.8.
(i) Note that in the covering we fix the $n$-th direction of $\Omega$, but allow rotating $\Omega$ in the first n1n-1 directions. This requires the two side face of $\Omega$ in the $n$-th direction both tile ${\mathbb{R}}^{n-1}$, moreover, the two side face have the same surface measure.
(ii) Let $\Omega$ be a strip-tiling domain that can cover $\mathbb{R}^{n-1}\times(0,w(\Omega))$. Let us denote by $S_{n}(\Omega)$ one of the two side faces of $\Omega$ in the $n$-th direction in what follows. Then the measure of $\Omega$ equals $|S_{n}(\Omega)|\times w(\Omega)$.
@@
Suppose that $\Omega$ is a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$. Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction. Assume that $\Omega_{0}$ is a piecewise smooth domain contained in $\Omega$. There exists $C_{P}=C(\Omega_{0},n)$ such that if $|S_{n}(\Omega)|\geq C_{P}$ then the Pólya conjecture holds on $\Omega\setminus\overline{\Omega}_{0}$.
Our first observation is that eigenvalues on strip-tiling domains satisfying better estimates than Pólya conjectured.
Theorem 1.10.
Let $\Omega$ be a strip-tiling domain that can tile ${\mathbb{R}}^{n-1}\times(0,w(\Omega))$, $n\geq 2$. Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction. Then it holds that for all $k\geq 1$ that
$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-C_{1}(n-1)|S_{n}(\Omega)|\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}},$
where $C_{1}(n)$ is as in (1.11).
[…]
Corollary 1.11.
[…]
$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{1}(n-1)}{K}|S_{n}(\tilde{\Omega})|\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}.$
The above result applies especially to any triangles in the plane.
Corollary 1.12.
Let $\Omega$ be a triangle in ${\mathbb{R}}^{2}$ with vertices at $a,b,c$. Then it holds for all $k\geq 1$ that
$k$ $\leq\frac{|\Omega|}{4\pi}\lambda_{k}(\Omega)-\frac{\max\{|ab|,|bc|,|bc|\}}{12\pi}\lambda_{k}(\Omega)^{\frac{1}{2}}.$
Remark 1.13.
Note that previously, it was only known that on an arbitrary triangle $\Omega$,
$k$ $\leq\frac{|\Omega|}{4\pi}\lambda_{k}(\Omega),\ \forall\ k\geq 1,$
which is due to Pólya [35]. In special cases of equilateral triangle, see Pinsky [33].
@@
Let $\Omega$ be a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$. Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{\mathbb{R}}}^{n}$}, that is contained in $\Omega$. If $|S_{n}(\Omega)|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then the Pólya conjecture holds on $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where
$C_{2}(n)=\begin{cases}2\sqrt{2}\pi,&n=1,\\ 9\sqrt{3},&n=2,\\ \frac{6(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{\mathscr{B}(\frac{n}{2}-1,2)}&n=3,4,\\ \frac{16(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{3\mathscr{B}(\frac{n}{2}-1,2)}\left(\frac{256}{247}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$ (1.12)
@@
Let $\Omega$ be a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$. Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction. Let $M\geq 2$ and $\Omega_{0}$ $M$-tile $\Omega$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{\mathbb{R}}}^{n}$}, that is contained in $\Omega_{0}$. If $|S_{n}(\Omega)|\geq C_{2}(n-1)M\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $\Omega_{0}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$.
Note that $M$ isometries of the domain $\Omega_{0}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$ corresponds to the case that removing $M$ copies of $\{Q_{k}\}_{k\in{\mathbb{N}}}$ from $\Omega$, the conclusion is obvious.
@@
(i) Let us take $\ell(k)=2^{\lfloor k/2\rfloor}$ for $k\geq 2$, $\ell(1)$=1, and $\{Q_{1,1},\,\{Q_{k,j}\}_{k\geq 2,1\leq j\leq\ell(k)}\}$ be a sequence of cubes, where $\ell(Q_{k,j})=2^{1-k}$. Then $\mathscr{S}_{Q}\leq 2(2+\sqrt{2})$. So for a strip-tiling domain $\Omega$ with $|S_{n}(\Omega)|>8(1+\sqrt{2})\pi$, one can arbitrarily remove any domain, that is composed by all or some of these cubes $\{Q_{1,1},\,\{Q_{k,j}\}_{k\geq 2,1\leq j\leq\ell(k)}\}$ with non-overlapped interior, from $\Omega$ such that the remaining set satisfies Pólya’s conjecture. See Figure 2 for some examples.
(a) A strip-tiling domain with a diamond and a room-corridor removed